OpenStudy (anonymous):

PLEASE HELP ME!!... Choose the answer. Solve for K. k + (-45)= -18 A. -27 B. -63 C. 27 D. 63

OpenStudy (idealist10):

k+(-45)=-18 k-45=-18 k=-18+45 k=45-18 k=27 Answer: 27, C